# VoCo-LLaMA: Towards Vision Compression with Large Language Models

### v1

@@ 4.3 Results
It can be observed that as the number of VoCo tokens grows, the overall compression performance of the model shows an upward trend. Increasing the number of tokens within the range of fewer than 10 tokens results in a significant improvement in compression performance. Finally, when conducting 128 VoCo tokens, the model achieves an average compression performance retention rate of 97.7%, indicating that the performance loss due to compression is almost negligible when compressing into more than 100 tokens. Interestingly, we observe that when training with 128 VoCo tokens, the result on the SEED-Bench exceeds the initial performance upper bound of the model.
| Method | context | content | GQA | POPE | SQAI | VQAT |
| LLaMA-VID [12] | 1 | 16 | 58.2 | 83.1 | 67.4 | 50.8 |
| 1 | 4 | 56.2 | 83.5 | 68.7 | 49.1 | |
| 1 | 1 | 55.5 | 83.1 | 68.8 | 49.0 | |
| VoCo-LLaMA | 0 | 1 | 58.3 | 85.0 | 69.5 | 52.7 |
| VoCo-LLaMA + context [12] | 1 | 1 | 58.5 | 85.6 | 69.2 | 52.7 |
Method of Compression. We compare our method with LLaMA-VID on vision compression, specifically evaluating its full model that utilizes both context and content tokens. For a fair comparison, VoCo-LLaMA is trained under the exact same settings and applied the same visual encoder, EVA-G [55], in this experiment. As shown in Table 3, our method outperforms the previous approach across the board when using a single content compression token, even surpassing the performance of LLaMA-VID when it uses multiple context tokens. In particular, we could observe an improvement of 2.8 and 3.7 on GQA and VQAT benchmarks, respectively. Additionally, we examine the role of context tokens, which facilitate cross-modal interactions between vision tokens and text, in VoCo-LLaMA. The impact of context tokens is not as significant as demonstrated in LLaMA-VID, possibly as VoCo-LLaMA’s inherent design already facilitates cross-modal alignment.
Inference Efficiency. VoCo-LLaMA can bring significant efficiency and storage gains during the inference phase. Due to our model’s design, the representation of compressed image (i.e., transformer activations on top of VoCo tokens) can be stored and repeatedly utilized in the form of a KV cache. We conduct a comparative analysis of CUDA time, FLOPs, and KV Cache storage size during the inference process, and compare our method with the baseline method and the full caching method. The baseline method, as its name suggests, does not employ any caching strategy and directly encodes and infers images. In contrast, the full caching method stores the uncompressed transformer activations upon all vision tokens as KV caches. More specifically, we follow the approach of [56], storing the keys and values of each transformer layer. As displayed in Table 4, we conduct an inference efficiency analysis of three methods on a single NVIDIA A100-SXM4-40GB using identical lengths of text prompts and single-image inputs. Compared to the baseline model without caching strategy, VoCo-LLaMA achieves a significant reduction of 69.6% in CUDA time and 94.8% in FLOPs by caching compressed tokens with a KV Cache of length one. Relative to the full caching strategy, our method save 99.8% of cache storage while achieving lower CUDA time and FLOPs, demonstrating the inference efficiency gains brought by our vision compression approach. Please refer to Appendix B for further discussion and details for inference efficiency.
@@
| Method | Token | KV Cache | Storage | CUDA | FLOPs | |||
| Length | Memory (MB) | Time (ms) | (T) | |||||
@@
| Method | Visual | Res. | Image | MSVD-QA | MSRVTT-QA | ANet-QA | |||
| Encoder | Token | Acc | Score | Acc | Score | Acc | Score | ||
| Methods w/o Vision Compression | |||||||||
| FrozenBiLM [57] | CLIP-L | 224 | 256 | 32.3 | - | 16.8 | - | 24.7 | - |
| Video-LLaMA [35] | EVA-G | 224 | 256 | 51.6 | 2.5 | 29.6 | 1.8 | 12.4 | 1.1 |
| LLaMA-Adapter [58] | - | 224 | - | 54.9 | 3.1 | 43.8 | 2.7 | 34.2 | 2.7 |
| VideoChat [36] | - | 224 | - | 56.3 | 2.8 | 45.0 | 2.5 | 26.5 | 2.2 |
| Video-ChatGPT [39] | CLIP-L | 224 | 256 | 64.9 | 3.3 | 49.3 | 2.8 | 35.2 | 2.7 |
| BT-ADapter [37] | CLIP-L | - | - | 67.5 | 3.7 | 57.0 | 3.2 | 45.7 | 3.2 |
| Vista-LLaMA [59] | EVA-G | 224 | 256 | 65.3 | 3.6 | 60.5 | 3.3 | 48.3 | 3.3 |
| Chat-UniVi [34] | CLIP-L | 224 | 256 | 69.3 | 3.7 | 55.0 | 3.1 | 46.1 | 3.3 |
| Methods w/ Vision Compression | |||||||||
| LLaMA-VID [12] | EVA-G | 224 | 2 | 69.7 | 3.7 | 57.7 | 3.2 | 47.4 | 3.3 |
| VoCo-LLaMA | CLIP-L | 224 | 2 | 72.3 | 3.9 | 61.1 | 3.5 | 47.9 | 3.4 |
| 336 | 2 | 72.6 | 3.9 | 61.2 | 3.5 | 47.9 | 3.4 | ||
| 224 | 8 | 73.4 | 3.9 | 62.0 | 3.5 | 48.5 | 3.4 | ||
| 336 | 8 | 73.5 | 3.9 | 62.3 | 3.5 | 48.6 | 3.4 | ||
@@
Video Understanding. We further evaluate the performance of VoCo-LLaMA on three widely used video understanding benchmarks, reporting results for input image resolutions of 224 and 336, respectively. First, we discuss the video understanding methods that utilize vision compression. Ensuring fair comparison, we adopted the same compression ratio as previous method [12], compressing each video frame into 2 VoCo tokens for training and inference in the video understanding model. Our method consistently outperforms previous video compression methods across all three benchmarks. Specifically, on the MSVD-QA and MSRVTT-QA datasets, VoCo-LLaMA achieved accuracies of 72.3% and 61.1%, respectively, corresponding to absolute gains of 3.7% and 5.9% over the previous best methods. Moreover, our method achieves the highest scores of 3.9 and 3.5, respectively.
@@
In comparison to video understanding methods that do not employ vision compression, our approach, which represents each video frame with a mere 2 VoCo tokens, demonstrates strong competitiveness against methods that utilize 256 or more vision tokens per frame. To further explore the potential of VoCo-LLaMA, we opted to compress video frames into the number of VoCo tokens that exhibited the optimal compression performance within the 0 order of magnitude (i.e., 8 tokens). Notably, as we increase the number of tokens, our method effectively leverages additional visual information, resulting in improved performance. We also analyse the performance loss caused by the use of vision compression in video domain, as detailed in the Appendix D. In summary, VoCo-LLaMA capitalizes on the strengths of its efficient vision compression to achieve robust performance on video data, all while bypassing the intricate temporal modeling designs.
5 Limitations
VoCo-LLaMA presents a novel solution for vision compression, but several limitations remain. Our training paradigm, while effective in enabling the model to comprehend compression tokens, inadvertently diminishes its ability to understand uncompressed vision tokens. Moreover, our model struggles to effectively handle tokens with diverse fine-grained compression levels. Addressing these limitations could unlock the door to infinite-length input visual-language models.
@@ 6 Conclusion
In this paper, we propose VoCo-LLaMA, the first approach to compress visual information using large language models. By distilling the LLM’s understanding of vision tokens into a compact representation, we achieve efficient compression of vision tokens. Our method can compress hundreds of vision tokens into a single VoCo token, while minimizing visual information loss. VoCo-LLaMA significantly reduces cache storage and boosts efficiency during the inference stage. Moreover, our method exhibits exceptional performance in learning temporal understanding on video data with continuous training, showcasing its promising potential. In summary, our approach offers a promising solution for fully utilise the limited and valuable context window of VLMs, making them more scalable and efficient for various multi-modal applications.
Appendix A Additional Implement Details
A.1 Experiment Details
We elaborate on the training details and hyperparameter design of VoCo-LLaMA during the training process. Since our method only requires introducing VoCo tokens during the visual instruction tuning stage, the hyperparameters reported are specific to the fine-tuning stage. We followed [2, 7] to configure the majority of the model parameters. Table 6 provides a detailed list of hyperparameters involved in the model training phase. Under the aforementioned parameter settings, VoCo-LLaMA is fine-tuned on 8 NVIDIA A100-SXM4-40GB GPUs.
| Hyperparameter | Value |
| Overall batch size | 128 |
| Learning rate | 2e-5 |
| LR Scheduler | Cosine decay |
| DeepSpeed ZeRO Stage | ZeRO-3 |
| Optimizer | AdamW |
| Warmup ratio | 0.03 |
| Epoch | 2 |
| Weight decay | 0 |
| Precision | bf16 |
A.2 Instruction Tuning Details
In this subsection, we delve into the design of instructions and special tokens for VoCo-LLaMA during the instruction tuning stage, illustrating the details with concrete examples. As shown in LABEL:lst:app_list_1, by leveraging the pre-trained language model Vicuna[11] as the foundation for our method, we ensured a consistent system prompt and dialogue design to prevent the model from learning or adapting to new instruction formats during fine-tuning.
Furthermore, the VoCo-LLaMA architecture necessitates organizing the input sequence in a specific order, i.e., vision tokens, VoCo tokens, and text tokens. However, demonstrated in LABEL:lst:app_list_1 [Common VLM], the standard "Conversation" and "Input" formats reveal that the special token <image> and its corresponding encoded <Vision Tokens> are situated in the middle of the dialogue sequence. To address this, we made modifications to align them in [VoCo-LLaMA]. We relocated the special token <image> to the forefront of the input prompt and appended the <VoCo Tokens> for compression. Concurrently, we refined the system prompt to accommodate the adjusted vision token position, thereby enabling the LLM to better comprehend and model the location of vision token.
Appendix B Inference Efficiency Details
B.1 CUDA Time
We present a comparative efficiency analysis of VoCo-LLaMA, the baseline and full-caching strategies during the inference phase in Table 4. Here, we provide additional details on the CUDA time measurement during the inference phase. We primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), kv cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. For the baseline model, which is a regular visual language model, the inference process requires additional visual information encoding, and, due to the lack of caching, the model must compute the transformer activations and generate text on all tokens (uncompressed visual and textual tokens) during inference.
Compared to the Full Caching strategy, the improvement in CUDA time achieved by VoCo-LLaMA can be attributed to two primary factors: Firstly, the time required to load the kv cache is significantly reduced, owing to the smaller storage size and shorter read time resulting from the 576 compression multiplier achieved by our approach. Secondly, the transformers forward time is also improved, although this is largely dependent on the length of the new input token (i.e., text token), which is identical in both cases. However, VoCo-LLaMA’s shorter prefix enables more efficient computation, leading to faster processing times. Notably, since both approaches cache transformer activations on top of vision tokens or compressed tokens, the image encoding duration is not a contributing factor to the overall CUDA time.
B.2 Storage Memory and FLOPs
The storage memory reported in Table 4 corresponds to the full transformer activations on top of the vision tokens or compressed VoCo tokens. This also refers to the size of the kv cache loaded during the inference process. Notably, the storage precision is bf16, which is consistent with the precision used in our training process.
Appendix C Implementation Code
LABEL:lst:app_list_2 presents the concrete implementation of introducing VoCo tokens and the attention mask mechanism in VoCo-LLaMA, corresponding to Equations 2 and 3. To facilitate reader understanding, we provide a brief code snippet implemented using PyTorch, accompanied by annotations.
Appendix D Video Compression Analysis
We have analyzed the vision compression performance in Table 1, and in this subsection, we further analyze and discuss the video compression performance. The setup of the boundary model is similar to the one described in Section 4.2, with the exception that the vision token input from a single image is replaced by a sequence of vision tokens from multiple video frames. Considering methods such as [34], our final setup samples 32 frames of a video during both training and inference phases. To ensure that the context window of the boundary model can accommodate the input length of the video frame sequences, we employ a 224224 resolution input for the frames and enlarged the context window of its language model by a factor of four using [60].
As shown in Table 7, VoCo-LLaMA achieves nearly lossless compression of video frames at a rate of 256, achieving an average compression multiplier of 97.4% on three widely-used video QA benchmarks. As the input image count increases, the vision compression operation of VoCo-LLaMA can reduce the total length of the vision token, mitigate the loss of information at a distance, and bring further gains to visual understanding.
| Method | Token | MSVD-QA | MSRVTT-QA | ANet-QA | Avg. | |||
| Acc | Score | Acc | Score | Acc | Score | |||
| Upper Bound | 256 | 68.8 | 3.7 | 57.3 | 3.2 | 46.5 | 3.3 | - |
| 100% | 100% | 100% | 100% | 100% | 100% | 100% | ||
| VoCo-LLaMA | 1 | 68.1 | 3.7 | 56.6 | 3.1 | 45.9 | 3.3 | - |
| 97.2% | 100% | 97.5% | 92.3% | 97.1% | 100% | 97.4% | ||
| Lower Bound | 1 | 43.5 | 2.2 | 29.4 | 1.9 | 25.9 | 2.2 | - |
| 0% | 0% | 0% | 0% | 0% | 0% | 0% | ||
Appendix E Visualized Examples of Output
Figure 5 illustrates the visualized examples of VoCo-LLaMA for vision compression. We compare our approach, which employs a single VoCo token, with the Upper Bound model described in Section 4.2 (denoted as U.B. throughout the figure). Correct responses are highlighted in green, while incorrect responses are marked in red. We further show some of the failure cases. For questions that require attention to finer details of the input image, such as the third question in the second case, VoCo-LLaMA sometimes produces incorrect descriptions. Notably, when the VoCo token count increases to 8, these erroneous outputs are significantly improved, as the compression retention rate rises.
Appendix F Broader Impacts
Our research relates to the Large Language Models and Large Vision Language Models. Although the training data we employ is public and compliant, there are potential negative societal impacts to consider. One possible concern is the generation of unintended or malicious outputs, which could be used to spread disinformation. To mitigate these risks, we propose that future work focus on developing mechanisms for monitoring and detecting misuse, as well as implementing gated release strategies for models to ensure they are used responsibly.

### v2

@@ 4.3 Results
Number of VoCo tokens. We evaluate the impact of the number of VoCo tokens on vision compression performance. Tab. 2 illustrates the trend of compression performance retention as the number of VoCo tokens varies, where the green and red lines represent the upper and lower bounds of compression performance, respectively. We adopted the same training settings and data as in the main experiments. It can be observed that as the number of VoCo tokens grows, the overall compression performance of the model shows an upward trend. Increasing the number of tokens within the range of fewer than 10 tokens results in a significant improvement in compression performance. Finally, when conducting 128 VoCo tokens, the model achieves an average compression performance retention rate of 97.7%, indicating that the performance loss due to compression is almost negligible when compressing into more than 100 tokens. Interestingly, we observe that when training with 128 VoCo tokens, the result on the SEED-Bench exceeds the performance upper bound model.
@@
| Method | Token | KV Cache Length | Storage Memory (MB) | CUDA Time (ms) | FLOPs (T) | |||
@@
| Method | Visual Encoder | LLM | Res. | Image Token | MSVD-QA | MSRVTT-QA | ActivityNet-QA | |||
| Acc | Score | Acc | Score | Acc | Score | |||||
| Methods w/o Vision Compression | ||||||||||
| FrozenBiLM [55] | CLIP-L | DeVERTa-V2 | 224 | 256 | 32.3 | - | 16.8 | - | 24.7 | - |
| Video-LLaMA [58] | EVA-G | Vicuna-7B | 224 | 256 | 51.6 | 2.5 | 29.6 | 1.8 | 12.4 | 1.1 |
| VideoChat [26] | - | Vicuna-7B | 224 | - | 56.3 | 2.8 | 45.0 | 2.5 | 26.5 | 2.2 |
| Video-ChatGPT [39] | CLIP-L | Vicuna-7B | 224 | 256 | 64.9 | 3.3 | 49.3 | 2.8 | 35.2 | 2.7 |
| BT-ADapter [34] | CLIP-L | Vicuna-7B | - | - | 67.5 | 3.7 | 57.0 | 3.2 | 45.7 | 3.2 |
| Vista-LLaMA [38] | EVA-G | Vicuna-7B | 224 | 256 | 65.3 | 3.6 | 60.5 | 3.3 | 48.3 | 3.3 |
| Chat-UniVi [23] | CLIP-L | Vicuna-7B | 224 | 256 | 69.3 | 3.7 | 55.0 | 3.1 | 46.1 | 3.3 |
| Methods w/ Vision Compression | ||||||||||
| LLaMA-VID [28] | EVA-G | Vicuna-7B | 224 | 2 | 69.7 | 3.7 | 57.7 | 3.2 | 47.4 | 3.3 |
| VoCo-LLaMA | CLIP-L | Vicuna-7B | 224 | 2 | 72.3 | 3.9 | 61.1 | 3.5 | 47.9 | 3.4 |
| 336 | 2 | 72.6 | 3.9 | 61.2 | 3.5 | 47.9 | 3.4 | |||
| 224 | 8 | 73.4 | 3.9 | 62.0 | 3.5 | 48.5 | 3.4 | |||
| 336 | 8 | 73.5 | 3.9 | 62.3 | 3.5 | 48.6 | 3.4 | |||
Method of Compression. We compare our method with LLaMA-VID on vision compression, specifically evaluating its full model that utilizes both context and content tokens. For a fair comparison, VoCo-LLaMA is trained under the exact same settings and applied the same visual encoder, EVA-G [14], in this experiment. As shown in Tab. 3, our method outperforms the previous approach when using a single content compression token, even surpassing the performance of LLaMA-VID when it uses multiple context tokens. In particular, we could observe an improvement of 2.8 and 3.7 on GQA and VQAT benchmarks, respectively.
Adaptability of VoCo Number. To assess the model’s adaptability to varying numbers of compression tokens, we trained the model with a fixed number of tokens and evaluated its performance with different token numbers. As demonstrated in Tab. 4, we conducted experiments by fixing the number of VoCo tokens (32) during training and varying the number of tokens during inference. Our method achieves better performance with an increasing number of compressed tokens, without requiring specialized training for elastic compressed token.
Results on fine-grained tasks. We analyze the extent of loss of fine-grained visual information after high-magnification compression of vision tokens in our approach. Here, we apply our method to [8] which is a cleanly structured MLLM trained on fine-grained task data such as REG, REC, and PointQA. As shown in Tab. 5 and Tab. 6, when compressing vision tokens to 1 VoCo token, our method maintained an impressive average compression retention rate of 79.9% and 81.6% for REC and REG tasks, respectively. Furthermore, by increasing the number of VoCo tokens to 8, we observed a significant improvement in the average compression retention rate. We observe that VoCo-LLaMA achieves similar compression retention rate to other benchmarks on fine-grained tasks, mainly because the Lower Bound model incurs more information loss on fine-grained tasks. Please refer to the supplementary material for additional fine-grained benchmarks, including VisWiz, OCRBench and others.
Inference Efficiency. We discuss the inference efficiency under the scenarios that images are cached as discussed in Sec. 3.3. Due to our model’s design, the representation of compressed image (i.e., transformer activations on top of VoCo tokens) can be stored and repeatedly utilized in the form of a KV cache. We conduct a comparative analysis of CUDA time, FLOPs, and KV Cache storage size during the inference process, and compare our method with the baseline method and the full caching method. The baseline method, as its name suggests, does not employ any caching strategy and directly encodes and infers images. In contrast, the full caching method stores the uncompressed Transformer activations upon all vision tokens as KV caches. More specifically, we follow the approach of [42], storing the keys and values of each Transformer layer. As displayed in Tab. 7, we conduct an inference efficiency analysis on a single NVIDIA A100 using identical lengths of text prompts and single-image inputs. Compared to the baseline model without caching strategy, VoCo-LLaMA achieves a significant reduction of 69.6% in CUDA time and 94.8% in FLOPs. Relative to the full caching strategy, our method save 99.8% of cache storage while achieving lower CUDA time and FLOPs, demonstrating the inference efficiency gains brought by our approach. Please refer to the supplementary material for further discussion and details for inference efficiency.
@@
Video Understanding. We further evaluate the performance of VoCo-LLaMA on three widely used video understanding benchmarks, reporting results for input image resolutions of 224 and 336, respectively. First, we discuss the video understanding methods that utilize vision compression. Ensuring fair comparison, we adopted the same compression ratio as previous method [28], compressing each video frame into 2 VoCo tokens for training and inference. Our method consistently outperforms previous video compression methods across all three benchmarks. Specifically, on the MSVD-QA and MSRVTT-QA datasets, VoCo-LLaMA achieved accuracies of 72.3% and 61.1%, respectively, corresponding to absolute gains of 3.7% and 5.9% over the previous best methods. Moreover, our method achieves the highest scores of 3.9 and 3.5, respectively.
@@
In comparison to video understanding methods that do not employ vision compression, our approach, which represents each video frame with a mere 2 VoCo tokens, demonstrates strong competitiveness against methods that utilize 256 or more vision tokens per frame. To further explore the potential of VoCo-LLaMA, we opted to compress video frames into the number of VoCo tokens that exhibited the optimal compression performance within the 0 order of magnitude (i.e., 8 tokens). Notably, as we increase the number of tokens, our method effectively leverages additional visual information. We also analyze the performance loss caused by vision compression and evaluate on other video QA benchmarks, as detailed in the supplementary material.
@@ 5 Conclusion
In this paper, we propose VoCo-LLaMA, the first approach to compress visual information using LLMs. By distilling the LLMs’ understanding of vision tokens into a compact representation, our method can compress hundreds of vision tokens into a single VoCo token, while minimizing information loss. VoCo-LLaMA significantly reduces cache storage and boosts efficiency during the inference stage. Moreover, our method exhibits promising performance in learning temporal understanding on video data with continuous training. In summary, our approach offers a promising solution for fully utilise the limited context window of VLMs, making them more efficient for multi-modal applications.